OpenStudy (anonymous):

Rewrite sin^4(x) so that it involves only the first power of cosine.

OpenStudy (anonymous):

@Spacelimbus can you help maybe?

OpenStudy (anonymous):

I am not quite sure about the question itself, do you know what you're currently studying? Double angular formulas for trigonometric identities? Or have you heard of Demoivre's theorem?

OpenStudy (anonymous):

i am doing trig identities, with double angle identities yeah

OpenStudy (anonymous):

and have you had an introduction to complex numbers? In special Demoivre?

OpenStudy (anonymous):

no i've not heard of demoivre

OpenStudy (anonymous):

Because, although I understand the nature of this problem, all I could do would give you some formulas which you would have to take as granted, there is not much more to it unfortunately :( On the other hand, if you know a bit about complex numbers then you could derive a formula yourself using \[\large \text{cis}(\alpha)= \cos \alpha + i \sin \alpha \implies \text{cis}(\alpha)^4= \text{cis}(4 \alpha)= \cos(4 \alpha) + i \sin(4 \alpha) \]

OpenStudy (anonymous):

is there any way to do it with just the double angle or squared identities ?

OpenStudy (anonymous):

Give me a moment to think, maybe I will manage to come up with something, you know for instance that: \[\large \sin (x) = \frac{1}{2}(1- \cos(2x)) \]

OpenStudy (anonymous):

Sorry it's supposed to be \(\sin ^2 (x)\)

OpenStudy (anonymous):

so you can square this identity and then try to come up with a double angular formula on the right hand side of it,

OpenStudy (anonymous):

seems to be at least one step into the correct direction, but it might still get tedious since we want to make the \( \cos ^2(2x)\) term after quadrating vanish. Can you follow so far?

OpenStudy (anonymous):

\[\large \sin ^4(x)= \left( \frac{1}{2}(1- \cos(2x) \right)^2= \frac{1}{4}(1 - 2 \cos(2x) + \cos^2(2x)) \]

OpenStudy (anonymous):

yes i follow that

OpenStudy (anonymous):

good, so for the very right identity, name \[\large \cos^2(2x)=1- \sin ^2(2x) \] right? And now apply the double angular theorem to \(\sin^2 (2x)\) again as above and you're done.